OpenStudy (anonymous):

Mr. Joe went to a store and bought 5 things for $65. He hat to pay 8.25% sales tax and had a discount coupon of 15% off. He gave the cashier a $100 bill. What was his change? i got $40.19

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

8.25% of sales tax that is 8.25% of 65 +65= 70.3625 and then 15% discount 70 -15% of 70.3625 = 70 -10.55= $59.45

OpenStudy (anonymous):

Hmm. You're sure?

OpenStudy (ghazi):

yep :)

OpenStudy (anonymous):

Okay. thanks:)

OpenStudy (anonymous):

Let me erase, then.

OpenStudy (ghazi):

its alright :)

OpenStudy (anonymous):

I did 65+5.36 (tax) = 70.36 -15% = 59.80 100-59.80 = 40.19

OpenStudy (ghazi):

you are absolutely right :)

OpenStudy (anonymous):

:o really?

OpenStudy (ghazi):

i forgot the last step that is 100-59.80

OpenStudy (anonymous):

Oh, okay. Tnx for checking, young man!

OpenStudy (ghazi):

you're welcome old lady :)

OpenStudy (anonymous):

:D